### v1

@@ 3.2.2 Defense with Gaussion Noise Injection
According to Eq. 5-11, the gradient of $W_{n}$ is closely related to the value of $h_{1}$. Thus, when Gaussion noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$ as $\tilde{h}_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations. In other words, injecting noise into the forward propagation process can protect the hidden states and its gradients transmitted during the training of the splitting LLMs.
3.3 Parallel Training
(a) Client-batch parallel training.
(b) Server-hierachical parallel training.
Figure 5: Multi-client Parallel Training of FL-LLaMA. In (a), the server receives forward-pass states from multiple clients simultaneously, concatenates them into a batch, and performs joint forward propagation. Then, the server splits hidden state $h_{B}$ and sends them back to different clients. In (b), each $client_{i}$ is paired with a corresponding $sub\_server_{i}$. After several training steps, the central server aggregates all sub-servers’ models using FedAvg and distributes the updated global server model back to each $sub\_server_{i}$.
In multi-clients settings, the FL-LLaMA framework accommodates three training paradigms: sequential training, client-batch parallelism and server-hierarchical parallelism. Each of them optimizes for different system configurations and resource availability. In sequential training, the server interacts with one client at a time in a round-robin manner, which is more suitable for scenarios with limited server capacity. In client-batch parallelism, the server waits for data from each client and then concatenate them in a single batch, which is appropriate for single machine with multiple GPUs server scenarios. In server-hierarchical parallelism, each client is assigned a separate sub-server for interaction, allowing multiple clients to train in true parallel, which is suitable for multi-machine, multi-GPU server environments.
3.3.1 Sequential Training
3.3.2 Client-Batch Parallelism
[…]
3.3.3 Server-hierarchical Parallelism
In server-hierarchical parallel setup, the central server spawns multiple sub-server model instances, each dedicated to a specific client, illustrated in Fig. 5(b). These instances run in parallel, allowing each client to perform local training simultaneously rather than waiting for others to complete. Typically after a fixed number of local updates, the central server synchronizes the training progress by aggregating model parameters across all sub-servers. This is usually done via weighted averaging of both server-side and client-side model parameters. This strategy enables significant improvements in training efficiency and scalability, especially in environments where the server has access to multiple GPUs or multi-threaded execution capabilities.
Table 1: Cloze questions and answers for the 8 SuperGLUE tasks
3.4 Inference Accelaration
@@
The inference time of LLMs are critically important for both enterprises and end-users. Low-latency inference minimizes user wait times, allowing for efficient and fast handling of complex tasks. Distributed architectures partition tensor operations, traditionally executed within GPUs and directly interacting with high-bandwidth memory (HBM), into separate computational units across multiple physical devices. Consequently, intermediate data must be transferred over networks, posing substantial latency challenges for the inference of LLMs. To enhance the efficiency and usability of distributed LLMs, our proposed FL-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
3.4.1 Attention Mask Compression
@@
FL-LLaMA transmits only the minimal information necessary for effective model training and inference. One significant improvement arises from reducing redundancy in attention mask transmissions. In masked self-attention mechanisms, the attention mask serves two primary purposes: (1) preventing decoders from prematurely accessing subsequent tokens during prediction, and (2) ensuring that padding tokens within input data do not interfere with attention computations involving valid tokens. Notably, the generation of attention masks across transformer blocks remains consistent throughout each training iteration and depends solely on dataset padding. Consequently, rather than transmitting the entire attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$, which incurs substantial bandwidth latency, the FL-LLaMA framework only sends essential metadata ($seq\_len$, $pad\_len$) indicating the locations of padding tokens versus valid tokens from the client to the server, thus significantly decreasing communication overhead.
3.4.2 Collaborative KV Cache Mechanism
@@
Since the model autoregressively generates the next token based on the tokens already generated, the speed of generating new tokens slows down as the input sequence length becomes excessively long (e.g., greater than 1000), due to the increasing number of K and V calculations. The application of key-value cache (KV cache) techniques in distributed inference further enhances inference speed. Generative LLMs produce outputs autoregressively, meaning that each generated token becomes part of the input sequence used to generate subsequent tokens. As user queries grow longer, repeatedly transmitting the complete context—comprising the initial query and previously generated tokens—from client to server can drastically increase inference latency. To mitigate this issue, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers. Since these states remain unchanged throughout token generation, newly generated tokens reuse cached states without requiring recomputation. Building upon this insight, FL-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step. This approach effectively minimizes data transmission, substantially improving inference efficiency.
3.5 Adaptive Partition
@@
Ege et al.[49] and Dario et al.[50] have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations. Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
@@
To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks in accordance with the demands of downstream tasks and available computational resources, thereby achieving an optimal balance between performance and efficiency. Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with Blocks_A=1,2,3) and the last blocks (with Blocks_C=1,2,3) of the model in Sec. 4.2.6. The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients. Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
4 Experiments
In order to demonstrate the performance of LLaMA within the federated split framework (FL-LLaMA), we conduct experiments using the same benchmarks as those used in LLaMA2-7B model [53]. Developed by Meta AI, the LLaMA2-7B model serves as a general-purpose language model with strong performance in natural language processing tasks, particularly when fine-tuned for specific domains. We employ this architecture to investigate whether FL-LLaMA can maintain comparable task performance to conventional centralized training, while addressing high delays and data privacy concerns through decentralized learning.
@@ 4 Experiments
Table 2: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FL-LLaMA and baselines.
[…]
4.1 Experimental Settings
@@
In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FL-LLaMA.
4.1.1 Dataset
@@
To evaluate model generalization across diverse NLP tasks, our experiments incorporates three distinct tasks:
  • 1.
    The SuperGLUE benchmark [58] is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, which span a broad range of linguistic challenges, from common sense reasoning to fine-grained semantic interpretation. Table 1 illustrates examples of cloze-style questions and corresponding answers derived from each task. (1)ReCoRD tests reading comprehension with common sense reasoning by resolving ambiguous entities in passages. (2)COPA assesses causal reasoning by choosing the more likely cause or effect for a given premise. (3)WSC focuses on pronoun disambiguation requiring deep contextual understanding. (4)RTE evaluates textual entailment by determining if one sentence follows, contradicts, or is unrelated to another. (5)BoolQ involves yes/no questions answered using evidence from context passages. (6)WiC tests word sense disambiguation by checking if a word has the same meaning in two sentences. (7)CB deals with entailment in complex sentences with nuanced reasoning. (8)MultiRC features multi-answer questions based on multi-sentence contexts, requiring integration of dispersed information. To unify these tasks under a single modeling framework, we adopt the formulation used in GLM [56], representing each instance as a blank infilling problem. Given a labeled pair $(x,y)$, the input $x$ is converted into a masked cloze question $q(x)$ using a placeholder token $[M]$, while the output $y$ serves as the target answer $a(y)$.
  • 2.
@@
    The CoQA dataset [59] is designed for the task of conversational question answering(QA), focusing on evaluating contextual coherence in multi-turn dialogues. It includes over 8,000 dialogues and more than 127,000 questions drawn from seven diverse domains. Answers are provided in free-form text, with nearly half of the questions requiring coreference resolution (e.g., pronoun disambiguation) and pragmatic reasoning. This setup closely mirrors real-world conversational settings, aiming to assess a model’s ability to understand text passages and respond to a series of interrelated questions.
  • 3.
    The XSum dataset [60] is designed for the task of abstractive summarization, specifically evaluating a model’s ability to perform extreme compression of salient information from news articles. It consists of BBC news articles paired with manually written summaries, covering a broad range of topics. A key challenge of XSum lies in its requirement for the model to generate a single-sentence summary for each article, placing significant demands on the model’s ability to abstract and condense information effectively.
4.1.2 Metrics
Since NLU tasks are reformulated as blank infilling problems, model performance is assessed based on the probability assigned to the ground-truth answer $a(y)$. For datasets such as RTE, BoolQ, WiC, CB, and MultiRC where the correct answer is often a single token, the evaluation score is derived from the logit corresponding to that token. Specifically, we define the probability as follows:
$p(y|x)=\frac{p(a(y)|q(x))}{\sum_{y^{{}^{\prime}}\in Y}p(a(y^{{}^{\prime}})|q(x))},$
where $Y$ represents the set of possible ground-truth labels.
In contrast, datasets like ReCoRD, COPA, and WSC require multi-token answers. In these cases, we evaluate model performance by summing the log-probabilities of each token in the predicted answer, computed as:
$s(y|x)=\sum_{t=1}^{|L_{y}|}\text{log}\text{P}(y_{t}|y_{1}...y_{t-1},x;\theta).$
@@ 4.1.2 Metrics
For conversational QA tasks, we utilize two metrics: Exact Match (EM) and F1 score to compare the model performance, which evaluate the accuracy of predicted answers against the ground truth [61, 62]. Exact Match (EM) measures the percentage of predictions that match the ground truth answers exactly, accounting for variations in punctuation and case. The F1 score is the harmonic mean of precision and recall, providing a more nuanced view of performance by rewarding partial overlaps between the predicted and reference answers.
@@
For summarization tasks, we adopt standard ROUGE metrics [63, 64, 65], i.e., ROUGE-1 and ROUGE-2, which are widely used to measure content overlap between generated summaries and reference texts. ROUGE-1 evaluates the unigram overlap, while ROUGE-2 measures the bigram overlap, providing insights into both content coverage and fluency.
4.1.3 Baselines
[…]
num. of clients 2 2 4 8 2 4 8 -
time(s) 689.2$\pm$5.7 630.0$\pm$1.8 542.2$\pm$3.5 488.6$\pm$0.4 348.2$\pm$3.8 173.53$\pm$3.0 86.4$\pm$0.6 381.9$\pm$2.4
4.1.4 Parameter Settings
@@
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation of the FL-LLaMA is based on the Flower [67] federated learning framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters.
4.2 Experimental Results
4.2.1 FL-LLaMA Performance
@@
In order to validate the effectiveness of FL-LLaMA, we compared the performance of FL-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 2. Experimental results show that FL-LLaMA and centralized LLaMA2-7B exhibit no significant performance differences, demonstrating the feasibility of the framework. Not only does FL-LLaMA perform comparably to centralized LLaMA2-7B, but the GPU memory consumption on the client side is also significantly reduced compared to centralized LLaMA2-7B. The comparison of their GPU memory usage is shown in Fig. 6. While centralized LLaMA2-7B occupies 27.9G of GPU memory, the FL-LLaMA client requires only 6G, representing a reduction of up to 82% in memory requirements. This reduction enables enterprises or departments that possess large amounts of high-quality private datasets but have limited computational resources to participate in federated training of LLMs.
@@
Figure 6: Computation load on client side of centralized LLaMA2-7B and FL-LLaMA on ReCoRD dataset.
4.2.2 Parallel Training
For communication cost analysis, we randomly selected 1,000 samples from the ReCoRD dataset. In our experimental setup, all training used a batch size of 1 and adopted FP16 training to improve computational efficiency. For LoRA, the decomposition rank $r$ is set to 8, and the scaling factor $\alpha$ is configured to 16, focusing specifically on the attention modules during fine-tuning. Additionally, the model is partitioned with Blocks_A=1 and Blocks_C=1. Each experiment is repeated five times, and we report the mean and standard deviation of the total communication time, as summarized in Tab. 3. The results show that sequential training takes the longest time, followed by client-batch training, and server hierarchical takes the shortest time. However, when the server can accommodate a larger batch size, increasing the number of clients in client-batch training effectively increases the batch size and thus reduces overall training time.

### v2

@@ Introduction
According to Eq. 1-7, the gradient of $W_{n}$ is closely related to the value of $h_{1}$. Thus, when Gaussian noise $\mathcal{N}(0,\delta)$ is injected into the hidden states $h_{1}$, perturbations are consequently introduced into gradients of parameters, thereby influencing subsequent gradient computations. In other words, injecting noise into the forward propagation process can protect the hidden states and their gradients transmitted during the training of the splitting LLMs.
Inference Accelaration
@@
The inference time of LLMs is critically important for efficient and fast handling of complex tasks. To enhance the efficiency and usability of distributed LLMs, our proposed FedSEA-LLaMA framework incorporates attention mask compression and collaborative KV cache mechanisms to minimize redundant data transmission.
Attention Mask Compression
@@
FedSEA-LLaMA transmits only the minimal information necessary for effective model training and inference. In particular, it reduces redundancy in attention mask transmissions compared to FL-GLM. Since the generation of attention masks remains consistent across transformer blocks within each training iteration and depends solely on dataset padding, transmitting full attention mask tensors $A_{mask}\in[0,-\infty]^{b\times seq\_len\times seq\_len}$ results in unnecessary communication overhead. Instead, FedSEA-LLaMA transmits only essential metadata (seq_lenseq\_len, pad_lenpad\_len) indicating padding locations, thereby significantly decreasing bandwidth usage.
Collaborative KV Cache Mechanism
@@
Since the model generates the next token autoregressively based on previously generated tokens, in distributed LLM setups, repeated transmission of the complete context from client to server can drastically increase inference latency as user queries grow longer. In centralized LLMs, the collaborative KV cache mechanism locally stores previously computed key and value states from the attention layers since these states remain unchanged throughout token generation. Building upon this insight, FedSEA-LLaMA caches the KV states locally at both the client and server sides, thereby restricting communication to the hidden state of the single newly-generated token at each inference step. This approach effectively minimizes data transmission, substantially improving inference efficiency.
Adaptive Partition
@@
Erdoğan, Küpçü, and Çiçek (2022) and Pasquini, Ateniese, and Bernaschi (2021) have pointed out that deeper model splits make it significantly harder for an attacker to recover information from the hidden states, as these data are produced through more complex computations. Furthermore, in real-world scenarios where the number of clients is large, the server may encounter computational resource bottlenecks, whereas clients are likely to have comparatively abundant computing resources.
@@
To flexibly accommodate specific task requirements and hardware constraints, we propose an adaptive splitting strategy that enables users to adjust the number of model input and output blocks, thereby achieving an optimal balance between performance and efficiency. Based on this strategy, we explore the performance of downstream tasks when the client holds the first blocks (with Blocks_A=1,2,3) and the last blocks (with Blocks_C=1,2,3) of the model. The results indicate that the depth of the model split does not significantly impact performance on downstream tasks, while affording a more flexible allocation of computational loads between the server and the clients. Furthermore, it corroborates the efficacy of our adaptive splitting strategy and delineates a refined configuration paradigm for deploying LLMs in federated learning environments.
Experiments
We conduct experiments on NLU, summarization generation, and conversational QA tasks to investigate whether FedSEA-LLaMA can maintain comparable performance to centralized training, while addressing high delays and data privacy concerns through decentralized learning. More experiments, including parallel training and training prompts, can be found in the Appendix.
Model ReCoRD COPA WSC RTE BoolQ WiC CB MultiRC CoQA Xsum F1/Acc. Acc. Acc. Acc. Acc. Acc. F1 F1a/EM F1/EM rouge-1/rouge-2 $\text{T5}_{\text{large}}$ (Raffel et al. 2020) 85.7/85.0 78.0 84.6 84.8 84.3 71.6 96.4 80.9/46.6 - 40.9/17.3 $\text{BART}_{\text{Large}}$ (Lewis et al. 2020) 88.3/87.8 60.0 65.4 84.5 84.3 69.0 90.5 81.8/48.0 - 45.1/22.3 $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) 89.6/89.0 82.0 83.7 87.7 84.7 71.2 98.7 82.4/50.1 - 45.5/23.5 ChatGLM-6B (Zeng et al. 2022) 80.2/78.7 85.0 71.2 81.6 83.4 71.0 85.7 78.2/45.6 - 37.6/12.5 LLaMA2-7B (Touvron et al. 2023) 81.3/79.5 75.0 68.3 73.3 83.8 70.9 85.7 82.3/50.1 88.6/80.9 45.9/26.1 FL-GLM (Zheng et al. 2024) 79.8/78.4 85.0 71.2 80.1 81.9 69.6 85.7 79.3/46.1 - 37.0/11.9 FedSEA-LLaMA 81.3/79.4 75.0 68.4 73.3 82.4 70.8 85.7 82.7/48.0 88.6/80.7 47.6/25.0
@@
Table 1: Performance Comparison on SupleGLUE, CoQA and Xsum datasets between FedSEA-LLaMA and baselines.
Experimental Settings
@@
In this section, we introduce some empirical settings, including datasets, evaluation metrics, baselines and parameter settings for FedSEA-LLaMA.
Dataset
@@
To evaluate model generalization across diverse NLP tasks, our experiments incorporate three distinct tasks:
  • The SuperGLUE benchmark (Wang et al. 2020) is designed for evaluating natural language understanding (NLU) tasks with eight distinct tasks, from common sense reasoning to fine-grained semantic interpretation.
@@
  • The CoQA dataset (Reddy, Chen, and Manning 2019) is designed for the task of conversational question answering(QA), featuring 8k+ dialogues (127k+ questions) across seven domains. Answers are free-form text, with nearly half of the questions requiring coreference resolution and pragmatic reasoning, simulating real-world dialogue coherence.
  • The XSum dataset (Narayan, Cohen, and Lapata 2018) evaluates abstractive summarization via extreme compression of BBC news articles into single-sentence summaries. It tests the model’s ability to abstract and condense salient information across diverse topics.
Metrics
We evaluate model performance using standard task-specific metrics. For NLU tasks such as COPA, WSC, etc, we report Accuracy, which measures the proportion of instances where the model’s predicted label exactly matches the ground-truth label.
@@
For conversational QA tasks, we use Exact Match (EM) and F1 score to evaluate model performance (Yatskar 2019; You et al. 2022). EM measures the percentage of predictions that exactly match the ground truth, accounting for punctuation and case, while F1 reflects the harmonic mean of precision and recall, rewarding partial overlaps.
@@
For summarization tasks, we use standard ROUGE metrics (Liu et al. 2021; Chen and Yang 2020; Fang et al. 2022), specifically ROUGE-1 and ROUGE-2, which assess unigram and bigram overlap between generated and reference summaries, capturing content coverage and fluency.
Baselines
We evaluate FedSEA-LLaMA based on the LLaMA2-7B model (Touvron et al. 2023), an open-source pre-trained model developed by Meta and FL-GLM. We additionally include several strong baselines: $\text{T5}_{\text{large}}$ (Raffel et al. 2020), $\text{BART}_{\text{Large}}$ (Lewis et al. 2020), $\text{GLM}_{\text{RoBERTa}}$ (Du et al. 2022) and ChatGLM-6B (GLM et al. 2024).
Parameter Settings
@@
The server is equipped with a Gigabit Ethernet card and utilizes multiple GPUs, including two NVIDIA A6000 and two NVIDIA L40. The implementation of the FedSEA-LLaMA is based on the Flower (Beutel et al. 2020) federated learning framework, which orchestrates the transfer of intermediate data between clients and the server. Due to limited computational resources, FP32 precision and a batch size of 1 are used only for the CoQA dataset. All other experiments are conducted with FP16 precision and a batch size of 2. The random seed is consistently set to 42. Fine-tuning is performed using distributed LoRA training, where $Blocks_{A}$, $Blocks_{B}$, and $Blocks_{C}$ each maintain their own LoRA adapters. For all LoRA fine-tuning experiments in this paper, the target modules are set to q_proj, v_proj, k_proj, o_proj.
Experimental Results
FedSEA-LLaMA Performance
@@
In order to validate the effectiveness of FedSEA-LLaMA, we compared the performance of FedSEA-LLaMA with centralized LLaMA2-7B on various tasks including NLU, summarization generation, and conversational QA, as shown in Table 1. Experimental results show that FedSEA-LLaMA achieves comparable performance to centralized LLaMA2-7B, demonstrating the feasibility of the framework while significantly reducing GPU memory consumption on the client side. The comparison of their GPU memory usage is shown in Figure 4. While centralized LLaMA2-7B occupies 28.2G of GPU memory, the FedSEA-LLaMA client requires only 3.4G, representing a reduction of up to 87.9% in memory requirements. Both models were finetuned on the Record dataset with a batch size of 2, using float16 precision and a LoRA rank of 8.
@@
Figure 4: Computation load on client side of centralized LLaMA2-7B and FedSEA-LLaMA on ReCoRD dataset.
(a) Loss Curves for Noise Scales
(Blocks_A=1, Blocks_C=1)
(b) Inference Speed Comparison
(w/ vs. w/o Attention Mask).
(c) Inference Speed Comparison
(w/ vs. w/o KV Cache).
Figure 5: (a)Different noise injection on MultiRC dataset. (b)Ablation study on attention mask. (c)Ablation study on KVcache.
Effects of Noise Scales
The corresponding experimental results are presented in Figure 5(a) and Table 2. Figure 5(a) demonstrates that when the number of blocks in $Blocks_{A}$ and $Blocks_{C}$ is held constant, the convergence speed of FedSEA-LLaMA decreases as the magnitude of injected noise increases. When the average absolute value of the noise is small (e.g., below 0.02), the loss curve remains close to that of the noise-free baseline. Table 2 indicates that under fixed noise levels, increasing the number of blocks in $Blocks_{A}$ consistently improves downstream performance in terms of both F1 and EM scores. Notably, when the noise amplitude reaches 0.1, only the single-block configuration of $Blocks_{A}$ shows a significant drop in downstream performance, whereas the three-block version of $Blocks_{A}$ demonstrates greater robustness.
Noise Scale Blocks_A = 1 Blocks_A = 2 Blocks_A = 3 0.0 77.48/40.82 77.82/39.24 77.89/41.34 0.02 77.89/38.09 77.87/39.66 78.61/42.39 0.05 74.57/29.17 77.07/38.41 78.95/41.03 0.1 59.11/8.18 72.45/23.71 75.68/34.10
Model Reverse Attack
To investigate the impact of split point selection and noise scale on the security of FedSEA-LLaMA, we evaluate the attack performance of the adversarial model under varying configurations of $Blocks_{A}$ depth (Blocks_A = 1, 2, 3) and noise (scale = 0, 0.02, 0.5). The attack model is trained via full-parameter fine-tuning with a learning rate of 2e-5. Reconstruction quality is evaluated using standard generation metrics BLEU-4 and ROUGE-2. FedBert reports a Rouge-2 of 7.053 and a Bleu-4 of 28.57. Experiment results of FedSEA-LLaMA is shown in Table  3. From the results, we can see that when the client holds only the embedding layer (similar to FedBERT), the attack model achieves relatively high BLEU and ROUGE scores. However, under the FedSEA-LLaMA framework, where the client also retains a portion of the LLMs block in addition to the embedding layer, all attack metrics approach zero. These findings demonstrate that FedSEA-LLaMA effectively mitigates model inversion attacks, thereby preserving the privacy of client-side data.
Noise scale Blocks_A=1 Blocks_A=2 Blocks_A=3 0.0 0.051/0.003 0.056/0.003 0.069/0.016 0.02 0.050/0.002 0.052/0.003 0.084/0.044 0.05 0.050/0.003 0.051/0.002 0.067/0.004
Table 3: Security analysis (Rouge-2 F1 / BlEU-4) of FedSEA-LLaMA on Record dataset.
Efficiency Optimization
Figure 5(b) shows a comparison of the time consumption between transferring the attention mask and transferring only the actual input length during distributed training. The bar chart represents the average time spent per token in FedSEA-LLaMA’s forward propagation, while the line chart represents the number of tokens generated per second. The context length is 500. It can be observed that transferring the compressed effective information of the attention mask plays a significant role in reducing the latency of distributed LLMs, with an average time reduction of 87.6%.
As shown in Figure 5(c), FedSEA-LLaMA demonstrates superior performance compared to FL-GLM in distributed inference, achieving a speed more than seven times faster. Without the KV cache, however, the inference speed decreases significantly as the query length increases, primarily due to the continuously growing number of K and V vectors. In contrast, with the KV cache technique, the client only needs to send the intermediate vector corresponding to a single token to the server, while reusing previously generated K and V vectors. This effectively avoids redundant computations. When the query length reaches 4000, the inference speed of FedSEA-LLaMA with KV cache is 8.2 times that without it. As query length increases further, this advantage continues to expand, greatly alleviating the computational burden and operational latency of distributed LLMs.
Adaptive Splitting
The choice of model partitioning has a significant impact on the computational burden at the client side. To evaluate whether FedSEA-LLaMA maintains robust performance under different partitioning strategies, we investigate its downstream task performance when clients hold different number of blocks in $Blocks_{A}$ and $Blocks_{C}$ (Blocks_A = 1, 2, 3 and Blocks_C = 1, 2, 3). This setting reflects realistic constraints where clients possess limited computational resources.
Blocks_A Blocks_C = 1 Blocks_C = 2 Blocks_C = 3 1 80.34/44.60 80.63/46.59 80.90/47.11 2 80.41/45.86 80.62/45.96 79.94/45.02 3 81.17/47.22 80.74/46.59 80.20/45.33
Table 4 shows F1a/EM scores across different partitioning schemes. Performance variations across partitions are minimal, indicating that FedSEA-LLaMA is robust to different split configurations. This flexibility allows resource-constrained clients to adjust block allocation based on their downstream task requirements and resource availability, without significant accuracy loss.
Conclusion
In this work, we proposed FedSEA-LLaMA, a novel federated split learning framework tailored for large language models. By securely offloading the majority of model computation to a central server while maintaining privacy-sensitive components on local clients, FedSEA-LLaMA addresses the core limitations of traditional federated split learning approaches. Through secure end-to-end propagation with noise injection, parallel training, inference acceleration, and adaptive partitioning tailored to downstream tasks and hardware, FedSEA-LLaMA demonstrates significant advances in security, efficiency, and adaptability. This federated split LLMs framework not only keeps users’ private data confined to their local devices but also significantly shortens training and inference time, making it better suited for handling the scale and complexity of LLMs. In future work, we aim to support more base models and deploy FedSEA-LLaMA framework in scenarios where data privacy is a critical concern.
Appendix A Noise Added to Hidden States